OpenStudy (anonymous):

a sail is in the form of a right triangle that is four times as high as it is wide. the sail is made from 32 square meter of material. what is its height??

OpenStudy (misssunshinexxoxo):

(1/2)(h/4)h = 32 h^2/8 = 32 h^2 = 32*8 = 256 h = sqrt(256) h = 16 Height is 16 meters.

OpenStudy (anonymous):

thank you

OpenStudy (misssunshinexxoxo):

Cheers